a. Company's float $60,000
b. Maximum payment
c. Maximum daily fee

Purple Feet Wine, Inc., receives an average of $20,000 in checks per day. The delay in clearing is typically three days. The current interest rate is .017 percent per day.

a. What is the company's float?
b. What is the most the company should be willing to pay today to eliminate its float entirely?
c. What is the highest daily fee the company should be willing to pay to eliminate its float entirely? (Round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

A) Company's float is given by the formula => average checks per day* number of days delay in clearing the check

= $20000*3= $6000

B) To eliminite its float the maximum amount Purple Fleet should be willing to pay is the Company's float which is $60000

C) Highest daily fee company should be willing to pay to eliminate its float is given by the formula company's float* intrest per day= 0.00017*60000=$ 10.20